# Transformer-Based Fault-Tolerant Control for Fixed-Wing UAVs Using Knowledge Distillation and In-Context Adaptation

### v1

@@ VI Conclusion
In this paper we have introduced a novel transformer-based approach for fault-tolerant control in fixed-wing UAVs, leveraging the transformer’s in-context learning capabilities to adapt in real-time to sudden changes in vehicle dynamics caused by structural and actuator failures. Unlike traditional FCSs, which struggle to maintain performance under substantial dynamic deviations, our method directly maps outer-loop reference values to control commands, bypassing the need for inner-loop controllers and additional fault detection mechanisms.
Through a teacher-student knowledge distillation framework, our approach trains a student policy on partial observations using knowledge transferred from a privileged expert agent. This enables the student to generalize across a range of complex failure scenarios, including actuator faults and structural damage, achieving high performance without requiring privileged environment information.
Experimental results demonstrate that our transformer-based control method not only outperforms industry-standard FCS and state-of-the-art reinforcement learning approaches but also remains resilient across challenging test cases. By effectively utilizing the transformer’s attention mechanism, our approach exhibits robust adaptability, maintaining control even in the face of severe aerodynamic and structural disruptions.
Future work will explore further enhancements to the model’s generalization capabilities by incorporating more complex training environments and testing on a wider array of fault conditions. Additionally, extending this transformer-based approach to other types of autonomous vehicles would help advance fault-tolerant control across various domains.

### v2

@@ VI Conclusion
In this paper, we introduced a transformer-based fault-tolerant flight control framework for fixed-wing UAVs that bypasses classical inner-loop controllers, drastically reducing the complexity of real-time adaptation to system failures. By coupling teacher-student knowledge distillation with a Decision Transformer trained in an offline RL setting, we demonstrated that partial observability can be effectively mitigated: the student learns robust policies solely from past state histories without explicit failure detection or online weight adaptation.
Our experiments reveal that the proposed controller achieves superior performance in both nominal and failure scenarios. In comparison with industry-standard PID-based FCS and other state-of-the-art RL agents, the transformer policy exhibited enhanced tracking accuracy and reduced crash rates, even under severe aerodynamic disruptions such as wing and tail damage. Through its attention mechanism, the Decision Transformer reallocates focus to post-failure observations, thereby adapting its control actions in context. This design also circumvents the need for real-time reparametrization or scheduled gain updates, offering a streamlined alternative to conventional fault-tolerant methods.
By incorporating Lyapunov-based constraints into the offline training process, the Decision Transformer is encouraged to explore actions that promote overall system stability, further improving safety in extreme conditions. The resulting lightweight model—comprising fewer than one million parameters—can be deployed efficiently on embedded processors, especially when combined with quantization. Our results thus highlight the potential of transformer architectures to unify real-time adaptation, fault tolerance, and computational feasibility in UAV control.
Future work will focus on hardware-in-the-loop and real-flight experiments to validate these promising simulation results under practical constraints such as limited onboard computation, sensor noise, and stringent flight-envelope requirements. Overall, this work serves as a step toward reliable, adaptive UAV operations that blend the robustness of classical control with the flexibility of modern deep learning methods, paving the way for safer and more versatile aerial robotic systems.
VII Acknowledgments
We thank Miguel Gómez-López and Rodney Rodríguez Robles for useful discussions on control theory.
S.L.C. acknowledges the grant PID2023-147790OB-I00 funded by MCIU/AEI/10.13039 /501100011033 /FEDER, UE. S.L.C. also acknowledges the MODELAIR project that have received funding from the European Union’s Horizon Europe research and innovation programme under the Marie Sklodowska-Curie grant agreement No. 101072559. The results of this publication reflect only the author’s view and do not necessarily reflect those of the European Union. The European Union can not be held responsible for them.
R.V. acknowledges financial support from ERC grant no.2021-CoG-101043998, DEEPCONTROL. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them.